# Autonomous Navigation and Station-Keeping on Near-Rectilinear Halo Orbits

### v1

@@ 4.3 $x$-axis Crossing Control as a Sequentially Linearized Minimization Problem
$\boldsymbol{\theta}^{\mathrm{EM}}(t_{f})=T^{\mathrm{Inr}}_{\mathrm{EM}}\left(\int_{t_{0}}^{t_{f}}\boldsymbol{f}(\tau,\boldsymbol{\theta}_{0})\mathrm{d}\tau+\begin{bmatrix}\Phi_{rv}(t_{f},t_{0})\\ \Phi_{vv}(t_{f},t_{0})\end{bmatrix}\boldsymbol{u}\right)=T^{\mathrm{Inr}}_{\mathrm{EM}}\boldsymbol{\theta}_{f,0}+B\boldsymbol{u},$ (37)
where $\boldsymbol{\theta}_{f,0}$ is the final state at $t_{f}$ if no control is executed at $t_{0}$, and
@@
Then, denoting $v_{x,0}^{\mathrm{EM}}(t_{f})=\left(T^{\mathrm{Inr}}_{\mathrm{EM}}\boldsymbol{\theta}_{f,0}\right)_{[4]}$, the linearized problem is given by
@@
Solving the minimization problem (39) is different from solving the root-finding problem (34) in two regards. First, the dynamics have been linearized using the STM in (37). Thus, the obtained solution $\boldsymbol{u}$ may not satisfy the constraint when propagated with the nonlinear IVP (32). In SK, the control that is sought is typically small in magnitude, and the linear approximation holds relatively well. Nevertheless, a few iterations (2 $\sim$ 3 in practice) solving (39) by updating $\boldsymbol{\theta}_{f,0}$ and $B$ may be necessary to incrementally obtain an aggregate control $\bar{\boldsymbol{u}}$ and satisfy the nonlinear constraint (33). Specifically, after the $i^{\mathrm{th}}$ iteration solving the linearized problem, $\boldsymbol{\theta}_{f,0}$ is replaced by $\boldsymbol{\theta}_{f,0}\leftarrow\boldsymbol{\theta}_{f,0}+\boldsymbol{u}^{(i)}$ where $\boldsymbol{u}^{(i)}$ is the solution from the $i^{\mathrm{th}}$ iteration, and $B$ is reconstructed using the updated initial state for the $i+1^{\mathrm{th}}$ iteration. Second, because this is a minimization of $\|\boldsymbol{u}\|_{2}$, the resulting control $\boldsymbol{u}$ is the minimum feasible $\boldsymbol{u}$ that simultaneously ensures the final $v_{x}^{\mathrm{EM}}$ to be within $v_{x,\mathrm{tol}}^{\prime}$ of the targeted value $\bar{v}_{x}^{\mathrm{EM}}$ according to the linear approximated dynamics (37).
Problem (39) is a SOCP, where the Euclidean norm in the objective may be formulated as a second-order cone constraint. Its optimal solution may be obtained analytically by observing that it is equivalent to determining the projection of the origin to the intersection of two half-spaces [22]. The control $\boldsymbol{u}$ satisfying (39b) is a vector $\boldsymbol{y}$ that belongs to the intersection of two half-spaces given by
@@
$\boldsymbol{\xi}_{1}^{T}$ $=B_{[4,:]},$ (41)
$\boldsymbol{\xi}_{2}^{T}$ $=-B_{[4,:]},$ (42)
$\eta_{1}$ $=v_{x,\mathrm{tol}}^{\prime}-v_{x,0}^{\mathrm{EM}}(t_{f})+\bar{v}_{x}^{\mathrm{EM}},$ (43)
$\eta_{2}$ $=v_{x,\mathrm{tol}}^{\prime}+v_{x,0}^{\mathrm{EM}}(t_{f})-\bar{v}_{x}^{\mathrm{EM}}.$ (44)
The expression for the projection of $\boldsymbol{y}$ onto $C$ is given by Proposition 28.19 in [37]
$P_{C}\boldsymbol{y}=\boldsymbol{y}-\nu_{1}\boldsymbol{\xi}_{1}-\nu_{2}\boldsymbol{\xi}_{2},$ (45)
@@
Since we seek to minimize $\|\boldsymbol{u}\|_{2}$, choosing $\boldsymbol{y}=\boldsymbol{0}_{3\times 1}$ in equation (45) yields the optimal solution $\boldsymbol{u}^{*}=P_{C}\boldsymbol{0}_{3\times 1}$ to problem (39).
In summary, the sequential linearization approach solves the NLP (31) by sequentially building the linearized problem (39) and computing its solution via (45). The overall procedure is termed the “sequentially linearized minimization problem” (SLMP), and is summarized in Algorithm 2, where the function sxform forms the transformation matrix at time $t_{f}$, and the function OriginProjectionCoefficients computes the coefficients $\nu_{1}$ and $\nu_{2}$ with $\boldsymbol{y}=\boldsymbol{0}_{3\times 1}$ using equation (46). The algorithm returns the aggregated control vector $\bar{\boldsymbol{u}}$, which is the summation of the solution $\boldsymbol{u}$ of each sequential instance of the linearized problem (39).
@@
$t_{0}$, $\boldsymbol{\theta}_{0}$, $t_{f}$, $\bar{v}_{x}^{\mathrm{EM}}$, $v_{x,\mathrm{tol}}$, $s$, $\mathrm{maxiter}$
@@
  $\eta_{1},\boldsymbol{\xi}_{1},\eta_{2},\boldsymbol{\xi}_{2}\leftarrow$ equations (41) $\sim$ (44)
@@ 4.4 Mean State Targeting via Unscented Transform
such that $\left|\mathbb{E}\left[\boldsymbol{\psi}^{\mathrm{B}}(t_{f})\right]-\bar{\boldsymbol{\psi}}^{\mathrm{B}}\right|\leq v_{x,\mathrm{tol}},$ (47b)
@@
where $\mathcal{Y}^{\ell}$ is given by
$\mathcal{Y}^{\ell}=T^{\mathrm{Inr}}_{\mathrm{EM}}\left(\mathcal{X}_{j-1\mid j-1}^{(\ell)}+\int_{t_{j-1}}^{t_{j}}\boldsymbol{f}\left(\tau,\mathcal{X}_{j-1\mid j-1}^{(\ell)}(\tau)\right)\mathrm{d}\tau\right),\quad\ell=0,\ldots,n,$ (49)
@@ 4.5 Trigger Condition and Targeting Tolerance Tuning
At each control action true anomaly, a trigger condition is used to determine whether a SK maneuver should be executed. Such a condition necessitates an indicator that quantifies the difference of the state estimate from the baseline. For the $x$-axis crossing control, a typical indicator is the difference in magnitude of $v_{x}^{\mathrm{EM}}$ at the $N^{\mathrm{th}}$ perilune with respect to the baseline. We define the threshold $v_{x,\mathrm{trig}}$ such that a control is triggered if $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|\geq v_{x,\mathrm{trig}}$. While both the DC and SLMP approaches use this dead-band $v_{x,\mathrm{trig}}$, the SLMP approach also requires appropriately tuning $v_{x,\mathrm{tol}}$.
When employing DC, the actual targeting violation $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|$ achieved is somewhat uncorrelated to $v_{x,\mathrm{tol}}$ due to the quadratic convergence nature; as such, the DC algorithm most often converges $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|\ll v_{x,\mathrm{tol}}$ Thus, as long as an appropriate $v_{x,\mathrm{trig}}$ is chosen, setting $v_{x,\mathrm{tol}}$ plays a relatively small role. The arguably arbitrary nature of the DC converging at different levels of $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|$ is an inherent hysteresis mechanism.
In contrast, the SLMP constrained minimization problem consistently aims at achieving $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|=v_{x,\mathrm{tol}}$ to minimize the control action. As such, $v_{x,\mathrm{tol}}$ plays a concrete role in determining how well the controller tracks the baseline. A tight $v_{x,\mathrm{tol}}$ results in expending more control effort to align the path closely to the baseline, while a relaxed $v_{x,\mathrm{tol}}$ results in the opposite behavior.
@@ 5 Filter Experiments
The first numerical experiment consists of running the synthetic image generation, processing it to generate position measurements, which are then fed to the recursive navigation filter. Results in this section do not include the station-keeping controller and focus on the performance of the filter alone. Without a controller, the trajectory naturally diverges after several revolutions due to the accumulation of numerical integration errors. As such, a simulation experiment without a controller can only last a limited number of revolutions. Meanwhile, the filter’s performance can be assessed within a few revolutions as the illumination conditions on an NRHO are almost periodic, and measurement qualities from one revolution to the next are nearly identical. Thus, in this section, for each considered filter configuration, 30 Monte-Carlo samples each lasting 10 revolutions are taken. The initial baseline epoch and state are given in Table 2. Across each Monte-Carlo run, the initial state estimate is computed by adding a random initial injection error of $3\sigma_{r}=10$  $\text{\,}\mathrm{k}\mathrm{m}$ and $\sigma_{v}=10$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$ to the baseline’s initial state. At each measurement instance, the camera is pointed toward the Moon based on the current state estimate; due to the estimation error, the boresight never exactly aligns with the center of the Moon.
We consider 12 measurement collection policies in total, summarized in Table 4. To abide by the typical operational requirement of a spacecraft, it is desirable to limit the number of measurements for navigational updates to a reasonably small number and low frequency. After initial investigations, we have found that taking a small number of images around the apolune region prior to the control is able to yield satisfactory navigation performance for station-keeping. Thus, we consider 3 and 4 measurement collection policies, where 2 or 3 measurements are taken before apolune in anticipation of a station-keeping maneuver, and one measurement is taken after such a maneuver and prior to the next perilune. For each case, the FOV is chosen so that an object with twice the apparent diameter of the Moon fits within the frame of the image at the closest range where a measurement is to be taken. For example, when collecting measurements at true anomalies of $140^{\circ}$, $150^{\circ}$, and $220^{\circ}$, the closest measurement occurs at a true anomaly of $140^{\circ}$ or $220^{\circ}$. While beyond the scope of this current work, one may choose to further optimize both the choice of the FOVs and the locations where the measurements are collected, but we focus on providing a sense of the trade-off that is in question.
@@
Canonical length $LU$, km 3000.0
Canonical time $TU$, s 2346.711856601253
@@ 5.2 EKF Performance
With the measurement configurations in Table 4, the EKF is able to track the state estimate without diverging/wrongfully converging. As an illustrative example, the EKF performance using a $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera is shown in Figure 6. While each run is randomized in terms of initial state error, error on the SRP coefficients of the true dynamics, and attitude error, the measurement itself is taken at locations within tens of kilometers difference from one Monte-Carlo run to another; thus, for any $i^{\mathrm{th}}$ measurement during a Monte-Carlo run, the analytical measurement covariance from equation (13) is nearly identical to that of the $i^{\mathrm{th}}$ measurement in any other Monte-Carlo run.
The difference in performance between each measurement configuration is visualized through the filter’s diagonal standard deviation history, shown against true anomaly ranges around apolune across the 10 revolutions in Figures 7 through 10. The red vertical lines indicate measurements, where the standard deviation sees a discrete drop due to the Kalman update. The standard deviation in velocity, which is of particular interest for station-keeping, follows a local minima that falls in the vicinity of apolune, but its exact location varies. Measurements with $f=300$  $\text{\,}\mathrm{m}\mathrm{m}$ or $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ result in lower $v_{z}$ standard deviation; $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ also sees slight improvement in standard deviations in $v_{x}$ and $v_{y}$, and is thus chosen as the measurement collection configuration for the GNC pipeline experiments in Section 7. While beyond the scope of this investigation, this experiment highlights the potential benefit of further optimizing and fine-tuning the camera FOV or measurement collection location and frequency for a specific orbit.
Figure 6: EKF performance with 3 measurements per revolution, $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera, for 30 Monte-Carlo runs; dark blue shaded region corresponds to 33-$\sigma$.
Figure 7: EKF filter standard deviation with 3 measurements per revolution, $f=300$  $\text{\,}\mathrm{m}\mathrm{m}$ camera. Red dashed lines indicate measurement collection.
Figure 8: EKF filter standard deviation with 3 measurements per revolution, $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera. Red dashed lines indicate measurement collection.
Figure 9: EKF filter standard deviation with 3 measurements per revolution, $f=450$  $\text{\,}\mathrm{m}\mathrm{m}$ camera. Red dashed lines indicate measurement collection.
Figure 10: EKF filter standard deviation with 3 measurements per revolution, $f=550$  $\text{\,}\mathrm{m}\mathrm{m}$ camera. Red dashed lines indicate measurement collection.

### v2

@@ 4.3 $x$-axis Crossing Control as a Sequentially Linearized Minimization Problem
$\boldsymbol{x}^{\mathrm{EM}}(t_{f})=T^{\mathrm{Inr}}_{\mathrm{EM}}\left(\int_{t_{0}}^{t_{f}}\boldsymbol{f}(\tau,\boldsymbol{x}_{0})\mathrm{d}\tau+\begin{bmatrix}\Phi_{rv}(t_{f},t_{0})\\ \Phi_{vv}(t_{f},t_{0})\end{bmatrix}\boldsymbol{u}\right)=T^{\mathrm{Inr}}_{\mathrm{EM}}\boldsymbol{x}_{f,0}+B\boldsymbol{u},$ (43)
where $\boldsymbol{x}_{f,0}$ is the final state at $t_{f}$ if no control is executed at $t_{0}$, and
@@
Then, denoting $v_{x,0}^{\mathrm{EM}}(t_{f})=\left(T^{\mathrm{Inr}}_{\mathrm{EM}}\boldsymbol{x}_{f,0}\right)_{[4]}$, the linearized problem is given by
@@
Solving the minimization problem (45) is different from solving the root-finding problem (40) in two regards. First, the dynamics have been linearized using the STM in (43). Thus, the obtained solution $\boldsymbol{u}$ may not satisfy the constraint when propagated with the nonlinear IVP (37). In SK, the control that is sought is typically small in magnitude, and the linear approximation holds relatively well. Nevertheless, a few iterations (2 $\sim$ 3 in practice) solving (45) by updating $\boldsymbol{x}_{f,0}$ and $B$ may be necessary to incrementally obtain an aggregate control $\bar{\boldsymbol{u}}$ and satisfy the nonlinear constraint (38). Specifically, after the $i^{\mathrm{th}}$ iteration solving the linearized problem, $\boldsymbol{x}_{f,0}$ is replaced by $\boldsymbol{x}_{f,0}\leftarrow\boldsymbol{x}_{f,0}+\boldsymbol{u}^{(i)}$ where $\boldsymbol{u}^{(i)}$ is the solution from the $i^{\mathrm{th}}$ iteration, and $B$ is reconstructed using the updated initial state for the $i+1^{\mathrm{th}}$ iteration. Second, because this is a minimization of $\|\boldsymbol{u}\|_{2}$, the resulting control $\boldsymbol{u}$ is the minimum feasible $\boldsymbol{u}$ that simultaneously ensures the final $v_{x}^{\mathrm{EM}}$ to be within $v_{x,\mathrm{tol}}^{\prime}$ of the targeted value $\bar{v}_{x}^{\mathrm{EM}}$ according to the linear approximated dynamics (43).
Problem (45) is a SOCP, where the Euclidean norm in the objective may be formulated as a second-order cone constraint. Its optimal solution may be obtained analytically by observing that it is equivalent to determining the projection of the origin to the intersection of two half-spaces [Elango2022SequentialLinear]. The control $\boldsymbol{u}$ satisfying (45b) is a vector $\boldsymbol{z}$ that belongs to the intersection of two half-spaces given by
@@
$\boldsymbol{\xi}_{1}^{T}$ $=B_{[4,:]},$ (47a)
$\boldsymbol{\xi}_{2}^{T}$ $=-B_{[4,:]},$ (47b)
$\eta_{1}$ $=v_{x,\mathrm{tol}}^{\prime}-v_{x,0}^{\mathrm{EM}}(t_{f})+\bar{v}_{x}^{\mathrm{EM}},$ (47c)
$\eta_{2}$ $=v_{x,\mathrm{tol}}^{\prime}+v_{x,0}^{\mathrm{EM}}(t_{f})-\bar{v}_{x}^{\mathrm{EM}}.$ (47d)
The expression for the projection of $\boldsymbol{z}$ onto $C$ is given by Proposition 28.19 in [BauschkeCombettes2010]
$P_{C}\boldsymbol{z}=\boldsymbol{z}-\nu_{1}\boldsymbol{\xi}_{1}-\nu_{2}\boldsymbol{\xi}_{2},$ (48)
@@
Since we seek to minimize $\|\boldsymbol{u}\|_{2}$, choosing $\boldsymbol{z}=\boldsymbol{0}_{3\times 1}$ in equation (48) yields the optimal solution $\boldsymbol{u}^{*}=P_{C}\boldsymbol{0}_{3\times 1}$ to problem (45).
In summary, the sequential linearization approach solves the NLP (36) by sequentially building the linearized problem (45) and computing its solution via (48). The overall procedure is termed the “sequentially linearized minimization problem” (SLMP), and is summarized in Algorithm 2, where the function sxform forms the transformation matrix at time $t_{f}$, and the function OriginProjectionCoefficients computes the coefficients $\nu_{1}$ and $\nu_{2}$ with $\boldsymbol{z}=\boldsymbol{0}_{3\times 1}$ using equation (49). The algorithm returns the aggregated control vector $\bar{\boldsymbol{u}}$, which is the summation of the solution $\boldsymbol{u}$ of each sequential instance of the linearized problem (45).
@@
$t_{0}$, $\boldsymbol{x}_{0}$, $t_{f}$, $\bar{v}_{x}^{\mathrm{EM}}$, $v_{x,\mathrm{tol}}$, $s$, $\mathrm{maxiter}$
@@
  $\eta_{1},\boldsymbol{\xi}_{1},\eta_{2},\boldsymbol{\xi}_{2}\leftarrow$ equations (47a) $\sim$ (47d)
@@ 4.4 Mean State Targeting via Unscented Transform
such that $\left|\mathbb{E}\left[\boldsymbol{\psi}^{\mathrm{B}}(t_{f})\right]-\bar{\boldsymbol{\psi}}^{\mathrm{B}}(\bar{t}_{f})\right|\leq v_{x,\mathrm{tol}}.$ (50b)
@@
where $\mathcal{Y}^{(\ell)}$ is the propagation of the $\ell^{\rm th}$ sigma point $\mathcal{X}_{j-1|j-1}^{(\ell)}$ transformed to $\mathcal{F}_{\rm EM}$,
$\mathcal{Y}^{(\ell)}=T^{\mathrm{Inr}}_{\mathrm{EM}}\mathcal{X}_{j\mid j-1}^{(\ell)}=T^{\mathrm{Inr}}_{\mathrm{EM}}\left(\mathcal{X}_{j-1\mid j-1}^{(\ell)}+\int_{t_{j-1}}^{t_{j}}\boldsymbol{f}\left(\tau,\mathcal{X}_{j-1\mid j-1}^{(\ell)}(\tau)\right)\mathrm{d}\tau\right),\quad\ell=0,\ldots,n,$ (52)
@@ 4.5 Trigger Condition and Targeting Tolerance Tuning
At each control action true anomaly, a trigger condition is used to determine whether a SK maneuver should be executed. Such a condition necessitates an indicator that quantifies the difference between the state estimate and the baseline. For the $x$-axis crossing control, a typical indicator is the difference in magnitude of $v_{x}^{\mathrm{EM}}$ at the $N^{\mathrm{th}}$ perilune with respect to the baseline. We define the threshold $v_{x,\mathrm{trig}}$ such that a control is triggered if $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|\geq v_{x,\mathrm{trig}}$. Setting $v_{x,\mathrm{trig}}>v_{x,\mathrm{tol}}$ builds in hysteresis aimed at avoiding chattering across revolutions, where
  1. (i)
    a control action is not engaged even if $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|>v_{x,\rm tol}$, until a larger tolerance $v_{x,\rm trig}$ is exceeded, and
  2. (ii)
    once engaged, the control action achieves $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|\leq v_{x,\rm tol}$, which is well below $v_{x,\rm trig}$.
Figure 6 illustrates hysteresis in the context of $x$-axis crossing control, where the horizontal axis corresponds to deviation in $v_{x}^{\rm EM}$, and the vertical axis indicates whether a control action is taken to decrease or increase $v_{x}^{\rm EM}$. A control action is taken within the shaded regions, and hysteresis corresponds to the blue shaded regions. With an appropriate tuning, the controller is effectively desensitized to the noise in the system, which is expected to lead to a reduction in cumulative $\Delta V$.
Figure 6: Hysteresis for $x$-axis crossing control
When employing DC, the actual targeting violation $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|$ achieved is somewhat uncorrelated to $v_{x,\mathrm{tol}}$ due to the quadratic convergence property of the Newton-Raphson scheme; as such, the DC algorithm most often converges $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|\ll v_{x,\mathrm{tol}}$. Thus, even with $v_{x,\mathrm{trig}}=v_{x,\mathrm{tol}}$, the control action will drive $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|$ below $v_{x,\rm trig}$, which in turns results in a similar effect as (ii) due to hysteresis.
In contrast, the SLMP constrained minimization problem consistently aims at achieving $|v_{x}^{\mathrm{EM}}-\bar{v}_{x}^{\mathrm{EM}}|=v_{x,\mathrm{tol}}$ to minimize the control action. As such, $v_{x,\mathrm{tol}}$ plays a concrete role in determining how well the controller tracks the baseline. For a given maneuver instance, a tighter $v_{x,\mathrm{tol}}$ results in expending more $\Delta V$ to align the path closely to the baseline, while a relaxed $v_{x,\mathrm{tol}}$ results in the opposite behavior. Meanwhile, the relationship between $v_{x,\mathrm{tol}}$ and the cumulative $\Delta V$ over multiple SK maneuvers is non-trivial. A tighter $v_{x,\mathrm{tol}}$ pushes the controlled trajectory closer to the reference, potentially leading to long-term, cumulative $\Delta V$ cost reduction over successive maneuver instances.
@@ 5 Filter Experiments
The first numerical experiment consists of running the synthetic image generation, processing it to generate position measurements, which are then fed to the recursive navigation filter. Results in this section do not include the station-keeping controller and focus on the performance of the filter alone. Without a controller, the trajectory naturally diverges after several revolutions due to the accumulation of numerical integration errors. As such, a simulation experiment without a controller can only last a limited number of revolutions. Meanwhile, the filter’s performance can be assessed within a few revolutions as the illumination conditions on an NRHO are almost periodic, and measurement qualities from one revolution to the next are nearly identical. Thus, in this section, for each considered filter configuration, 30 Monte-Carlo samples each lasting 10 revolutions are taken. The initial baseline epoch and state are given in Table 1. Across each Monte-Carlo run, the initial state estimate is computed by adding a random initial injection error of $3\text{-}\sigma_{r}=10$  $\text{\,}\mathrm{k}\mathrm{m}$ and $3\text{-}\sigma_{v}=10$  $\text{\,}\mathrm{c}\mathrm{m}\mathrm{/}\mathrm{s}$ to the baseline’s initial state. At each measurement instance, the camera is pointed toward the Moon based on the current state estimate; due to the estimation error, the boresight never exactly aligns with the center of the Moon. Simulation parameters are summarized in Table 2.
We consider 4 measurement collection policies in total, summarized in Table 3. To abide by the typical operational requirement of a spacecraft, it is desirable to limit the number of measurements for navigational updates to a reasonably small number and low frequency. After initial investigations, we have found that taking a small number of images around the apolune region prior to the control is able to yield satisfactory navigation performance for station-keeping. Thus, we consider 3-measurement collection policies, where 2 measurements are taken before apolune in anticipation of a station-keeping maneuver, and one measurement is taken after such a maneuver and prior to the next perilune. For each case, the FOV is chosen so that an object with twice the apparent diameter of the Moon fits within the frame of the image at the closest range where a measurement is to be taken. For example, when collecting measurements at true anomalies of $140^{\circ}$, $150^{\circ}$, and $220^{\circ}$, the closest measurement occurs at a true anomaly of $140^{\circ}$ or $220^{\circ}$. While beyond the scope of this current work, one may choose to further optimize both the choice of the FOVs and the locations where the measurements are collected, but we focus on providing a sense of the trade-off that is in question.
@@
Canonical length $\rm LU$, km 3000.0
Canonical time $\rm TU$, s 2346.711856601253
@@ 5.2 EKF Performance
With the measurement configurations in Table 3, the EKF is able to track the state estimate without diverging/wrongfully converging. As an illustrative example, the EKF performance using a $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ camera is shown in Figure 8. While each run is randomized in terms of initial state error, error on the SRP coefficients of the true dynamics, and attitude error, the measurement itself is taken at locations within tens of kilometers difference from one Monte-Carlo run to another; thus, for any $i^{\mathrm{th}}$ measurement during a Monte-Carlo run, the analytical measurement covariance from equation (15) is nearly identical to that of the $i^{\mathrm{th}}$ measurement in any other Monte-Carlo run. We also note the spike in velocity errors, along with the corresponding covariances, at perilune passes. The large error is due to the rapid rotation of the spacecraft’s velocity vector around perilune, leading to along-track error growth. Following perilune, the error quickly reduces due to the dynamics as well as the incoming measurement.
The difference in performance between each measurement configuration is visualized through the filter’s standard deviation history: Figure 9 shows the root mean squares of the EKF’s position and velocity standard deviation around apolune across 10 revolutions. The red vertical lines indicate measurements, where the standard deviation sees a discrete drop due to the Kalman update. The standard deviation in velocity, which is of particular interest for station-keeping, follows a local minima that falls in the vicinity of apolune where $\theta=180^{\circ}$, but its exact location varies. Measurements with $f=300$  $\text{\,}\mathrm{m}\mathrm{m}$ or $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ result in lower velocity standard deviation; $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$ also sees slight improvement in position standard deviations, and is thus chosen as the measurement collection configuration for the GN&C pipeline experiments in Section 7. While beyond the scope of this investigation, this experiment highlights the potential benefit of further optimizing and fine-tuning the camera FOV or measurement collection location and frequency for a specific orbit.
Figure 8: Monte-Carlo runs of EKF with 3 measurements per revolution using $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$
(a) $f=300$  $\text{\,}\mathrm{m}\mathrm{m}$, $\mathrm{FOV}\approx 18.92^{\circ}$
(b) $f=360$  $\text{\,}\mathrm{m}\mathrm{m}$, $\mathrm{FOV}\approx 15.81^{\circ}$
(c) $f=450$  $\text{\,}\mathrm{m}\mathrm{m}$, $\mathrm{FOV}\approx 12.68^{\circ}$
(d) $f=550$  $\text{\,}\mathrm{m}\mathrm{m}$, $\mathrm{FOV}\approx 10.39^{\circ}$
Figure 9: Root mean squares of EKF standard deviation with 3 measurements per revolution